OpenStudy (anonymous):

Factor y = x^3 - 4x^2 - 7x +10 if x=-2 is a zero.

OpenStudy (anonymous):

dont think too hard since \(-2\) is a zero, you know it looks like \[ x^3 - 4x^2 - 7x +10=(x+2)\times (\text{something})\] you can find the something by a) long division (lousy method) b) synthetic division (easy method) c) thinking (good method)

OpenStudy (anonymous):

It says to use synthetic, but I'm having trouble with it.